OpenStudy (trojanpoem):

Integrate:

OpenStudy (trojanpoem):

\[\int\limits_{}^{}\frac{ dx }{ \sqrt{e^x + lnx} }\]

OpenStudy (rsadhvika):

wolfram gave up http://www.wolframalpha.com/input/?i=Integrate%5B1%2Fsqrt(e%5Ex%2Blnx),%7Bx%7D%5D not sure if i want to attempt this..

OpenStudy (trojanpoem):

lol

OpenStudy (priyar):

is this your Q.? sure?

OpenStudy (trojanpoem):

It's my Professor's.

OpenStudy (kainui):

I'd be interested to see if anyone can make any kind of statement about this more general integral \[\int \frac{dx}{\sqrt{f(x) + f^{-1}(x)}}\]

OpenStudy (astrophysics):

.

OpenStudy (damonte212):

..

OpenStudy (kainui):

...